# GateLens: A Reasoning-Enhanced LLM Agent for Automotive Software Release Analytics

### v2

@@ Abstract
Ensuring reliable software release decisions is critical in safety-critical domains such as automotive manufacturing. Release validation relies on large tabular datasets, yet manual analysis is slow, costly, and error-prone. While Large Language Models (LLMs) offer promising automation potential, they face challenges in analytical reasoning, structured data handling, and ambiguity resolution. This paper introduces GateLens, an LLM-based system for analyzing tabular data in the automotive domain. GateLens translates natural language queries into Relational Algebra (RA) expressions and generates optimized Python code. Unlike traditional multi-agent or planning-based systems that can be slow, opaque, and costly to maintain, GateLens emphasizes speed, transparency, and reliability. Experimental results show that GateLens outperforms the existing Chain-of-Thought (CoT) + Self-Consistency (SC) based system on real-world datasets, particularly in handling complex and ambiguous queries. Ablation studies confirm the essential role of the RA layer. Industrial deployment shows over 80% reduction in analysis time while maintaining high accuracy across test result interpretation, impact assessment, and release candidate evaluation. GateLens operates effectively in zero-shot settings without requiring few-shot examples or agent orchestration. This work advances deployable LLM system design by identifying key architectural features—intermediate formal representations, execution efficiency, and low configuration overhead—crucial for safety-critical industrial applications.
Index Terms:
@@
I Introduction
@@
Validating software releases in the automotive industry is a multifaceted challenge, particularly for embedded software in safety-critical systems. Modern vehicles integrate numerous subsystems, making this process complex and resource-intensive. Each integration phase involves gating steps—critical checkpoints where tests verify compliance with predefined quality standards. Failures at these gates can ripple through the system, delaying dependent subsystems regardless of their individual quality. Release managers tasked with safeguarding quality must analyze vast quantities of test results and validation data. While essential for ensuring safety and reliability, this process is time-consuming and prone to human error in data interpretation.
@@
The software industry’s transition from manual to automated processes has entered a new era with the emergence of Large Language Models (LLMs) (Liu et al., 2024; Chang et al., 2024). Companies are increasingly integrating these AI agents into their workflows, seeking more cost-effective and optimized solutions for complex software engineering tasks (Leung and Murphy, 2023). However, direct application of LLMs to software release validation is hindered by limitations in interpretable reasoning and understanding of technical specifications (Marques, 2024; Austin et al., 2021).
To address these challenges, we introduce GateLens, a reasoning-enhanced LLM agent (Miehling et al., 2025) to support release validation in the automotive domain. GateLens integrates structured relational analysis with domain-specific expertise, leveraging a reasoning layer built on Relational Algebra (RA) to break down complex validation tasks into systematic analytical steps. GateLens simplifies three critical aspects of release validation:
@@
  • We design an architecture optimized for time- and safety-critical environments, minimizing LLM invocations while preserving reasoning depth. GateLens operates in a zero-shot setting, avoiding the need for few-shot examples or multi-agent coordination, which improves generalizability, execution speed, reduces maintenance overhead, and enhances transparency.
@@ II Background and Motivation
These limitations highlight the need for enhanced solutions combining domain expertise with advanced analytical reasoning capabilities. Such solutions must facilitate efficient decision-making while maintaining high safety and quality standards in automotive software development (Wang et al., 2024a). Effective intelligent assistants can transform the release decision process, enabling release managers to prioritize result interpretation and strategic decision-making over routine data analysis.
III Approach and Methodology
@@ III-B Core Components
III-B1 Query Interpreter
The query interpreter agent is responsible for converting user queries into formal RA expressions, providing a precise framework for analytical reasoning. Before initiating this translation, the agent consults the knowledge base, comprising the data schema and domain-specific context. The data schema provides a detailed understanding of the dataset, including its relational modeling, detailed field descriptions, and data types. Using this information, the agent verifies whether the query is relevant and within the scope of the dataset (Manik et al., 2021). This validation step ensures that only supported and meaningful queries are processed, improving both accuracy and efficiency. Once the query is confirmed to be in scope, the agent leverages the data schemas to interpret and decompose the query into formal RA expressions.
The agent’s primary function is to map natural language queries into formal RA expressions, enhancing LLM reasoning through structured decomposition (Khot et al., 2022). This approach extends traditional Chain-of-Thoughts (CoT) (Wei et al., 2022) reasoning by constraining the model to think within a formal system framework (Zhang et al., 2023a). Instead of generating free-form solutions, the agent must express analytics through a limited set of standard operations: selection, projection, union, set difference, cartesian product, and rename as basic operations, as well as derived operations such as join, intersection, and division and complemented by aggregation functions like average, minimum, maximum, sum, and count.
@@ III-B2 Coder
III-C Data Handling
A key architectural decision in GateLens is its indirect interaction with test data. Rather than exposing sensitive test data directly to LLM agents, which could raise privacy concerns (Boudewijn et al., 2023) and exceed input context limitations, the system operates on data schemas and relational models. This approach serves multiple critical purposes:
@@ 4. RQ4:
    How does the number of few-shot examples impact GateLens’s performance?
IV-A Experimental Setup
To address the research questions introduced in Section IV, we designed and conducted extensive experiments to evaluate the performance of GateLens.
@@ IV-A Experimental Setup
The second benchmark is derived from real-world user queries collected from production logs at our partner company. These queries were sourced from the historical logs of an agentic system that employed a well-established tabular data reasoning approach combining CoT (Wei et al., 2022) prompting with Self-Consistency (SC) (Wang et al., 2022). This system was used in production to support software release analytics, and the collected queries reflect a wide range of user roles, query types, and domain-specific requirements. While this system performs effectively in many scenarios, its limitations become apparent as the range of roles and users expands, leading to a significant diversification of queries. This broader query diversity exposes the system’s reliance on few-shot examples, making it less capable of handling highly complex, ambiguous, or ill-defined queries that require greater flexibility and adaptability. Nevertheless, this preliminary system played a critical role in data collection for GateLens by providing query logs used to develop and validate our approach. From these logs, we filtered out near-duplicates and selected 244 frequently repeated unique queries, which we then organized into eight functional categories based on their purposes.
In order to assess GateLens’s performance, we run experiments with two large language models (LLMs): GPT-4o, a leading commercial model, and Llama 3.1 70b, a recently released open-source model. We also benchmark GateLens against the CoT+SC agentic system currently used to support the company’s release decisions. Our comparative analysis is designed to quantify the improvements introduced by GateLens’s novel architecture.
To address the challenge of handling out-of-scope user queries during real-time interactions, GateLens incorporates an in-scope filtering mechanism as explained in Section III-B. This mechanism ensures that the system only attempts to process queries that fall within its scope, thereby improving reliability and reducing errors. Performance evaluation focused on two key aspects:
@@
Based on these definitions, we calculate Precision, Recall, and F1 scores to assess the performance of the system. Precision ensures that incorrect results are minimized, recall ensures relevant queries are addressed, and the F1 score balances the two to provide an overall assessment of system performance.
IV-B Performance in Addressing User Queries (RQ1)
TABLE I: Performance comparison across different models and difficulty levels on the first benchmark, which consists of 50 designed queries with annotated difficulty levels.
@@ IV-B Performance in Addressing User Queries (RQ1)
TABLE II: Performance comparison of GateLens and the CoT+SC system across different categories on the second benchmark, which consists of 244 real-world queries.
@@
We conducted experiments to compare the performance of GateLens across the two introduced benchmarks. The first benchmark, consisting of 50 queries categorized by difficulty levels, was used to evaluate and compare the performance of GateLens and CoT+SC. Both systems were tested using GPT-4o and Llama 3.1 70B as their underlying LLMs. The results are summarized in Table I.
@@
For the second benchmark, results in Table II show that GateLens (GPT-4o) and CoT+SC (GPT-4o) significantly outperformed GateLens with Llama 3.1 70B. This performance disparity is primarily due to the strict code generation requirements of the task, including table filtering, merging strategies, and key-value mapping operations, where GPT-4o demonstrated markedly superior capabilities.
@@
Across both datasets, GateLens with GPT-4o consistently delivered the best performance, achieving perfect accuracy across difficulty levels in the first benchmark and superior handling of diverse query categories in the second benchmark. This superior performance stems from the system’s ability to extract logical conditions from natural language queries and transform them into RA expressions, clarifying query logic, ensuring accurate, complete condition handling and minimizing errors and omissions in the generated code. These results demonstrate the framework’s enhancement of LLM performance on complex queries and improved real-world reliability
IV-C Robustness: Handling Out of Scope and Imprecise Queries (RQ2)
@@
IV-C1 Handling Out of Scope Queries
We compared GateLens with other models; the results can be found in Table III. The results demonstrate that GateLens with GPT-4o achieved the best performance, particularly in terms of precision, which is approximately 40% higher than other models, indicating GateLens’ ability to avoid generating incorrect results.
@@ IV-C1 Handling Out of Scope Queries
TABLE III: Model comparison for out-of-scope queries.
@@
IV-C2 Handling Imprecise Queries
@@
TABLE IV: Model comparison for imprecise queries.
@@
The results of these experiments are presented in Table IV. As shown, GateLens with GPT-4o demonstrates the best overall performance. In terms of precision, all methods performed relatively well, indicating that when results are generated, they are likely to be correct. However, our method significantly outperformed the others in recall, highlighting its ability to handle a larger portion of the imprecise queries. As a result, GateLens with GPT-4 achieved a substantially higher F1 score compared to other methods, demonstrating that it not only processes most queries but also produces accurate results for them.
@@
Overall, our approach demonstrated a strong ability to handle imprecise queries by maintaining high precision and significantly improving recall. This robustness ensures that the system is adaptable to variations and ambiguities in query formulations, allowing users to freely express their queries without being constrained by strict adherence to column header names. Such flexibility enhances the system’s practicality and applicability in real-world scenarios, making it a reliable and user-friendly tool for diverse interactions.
IV-D Effectiveness of the RA module (RQ3)

### v3

# arXiv:2503.21735v3[cs.SE] 01 Mar 2026

@@ Abstract
Ensuring reliable data-driven decisions is crucial in domains where analytical accuracy directly impacts safety, compliance, or operational outcomes. Decision support in such domains relies on large tabular datasets, where manual analysis is slow, costly, and error-prone. While Large Language Models (LLMs) offer promising automation potential, they face challenges in analytical reasoning, structured data handling, and ambiguity resolution. This paper introduces GateLens, an LLM-based architecture for reliable analysis of complex tabular data. Its key innovation is the use of Relational Algebra (RA) as a formal intermediate representation between natural-language reasoning and executable code, addressing the reasoning-to-code gap that can arise in direct generation approaches. In our automotive instantiation, GateLens translates natural language queries into RA expressions and generates optimized Python code. Unlike traditional multi-agent or planning-based systems that can be slow, opaque, and costly to maintain, GateLens emphasizes speed, transparency, and reliability. We validate the architecture in automotive software release analytics, where experimental results show that GateLens outperforms the existing Chain-of-Thought (CoT) + Self-Consistency (SC) based system on real-world datasets, particularly in handling complex and ambiguous queries. Ablation studies confirm the essential role of the RA layer. Industrial deployment demonstrates over 80% reduction in analysis time while maintaining high accuracy across domain-specific tasks. GateLens operates effectively in zero-shot settings without requiring few-shot examples or agent orchestration. This work advances deployable LLM system design by identifying key architectural features—intermediate formal representations, execution efficiency, and low configuration overhead—crucial for domain-specific analytical applications where accuracy, traceability, and stakeholder trust are paramount.
keywords:
@@
journal: Journal of Systems and Software
[…]
[label1]organization=Department of Computer Science and Engineering, Chalmers University of Technology, city=Gothenburg, country=Sweden
[…]
[label2]organization=Volvo Group, city=Gothenburg, country=Sweden
1 Introduction
@@
Reliable decision-making in data-intensive domains depends on the ability to accurately analyze large volumes of structured data. In sectors such as automotive manufacturing, healthcare, finance, and regulatory compliance, critical decisions hinge on interpreting tabular datasets that capture test results, operational metrics, or validation records. These datasets often pass through gating steps—critical checkpoints where predefined quality or compliance standards must be met. Failures at these gates can cascade through interconnected processes, delaying dependent workflows regardless of their individual quality. Analysts tasked with safeguarding decision quality must process vast quantities of data. While essential for ensuring accuracy and reliability, this process is time-consuming and prone to human error in data interpretation.
@@
The software industry’s transition from manual to automated processes has entered a new era with the emergence of Large Language Models (LLMs) [10.3390/app14031046, chang2024survey]. Companies are increasingly integrating these AI agents into their workflows, seeking more cost-effective and optimized solutions for complex software engineering tasks [10.1109/ase56229.2023.00035]. However, direct application of LLMs to structured data analysis for decision support is hindered by limitations in interpretable reasoning and understanding of technical specifications [10.3390/fi16060180, austin2021program].
Figure 1: Comparison of conventional Chain-of-Thought (CoT) reasoning and its extension using Relational Algebra (RA) statements. (Left) CoT employs unstructured, fused reasoning where multiple analytical concepts are blended together in informal thoughts without clear separation. Operations cannot be independently mapped to executable code snippets, making reasoning steps opaque and difficult to debug in isolation. (Right) RA-based reasoning adopts a structured, compositional approach where each relational algebra operation (filtering, joining, projecting) is a discrete, formally grounded step. Each operation is independently interpretable, reusable, and can be directly mapped to executable code. The color-coding illustrates how RA maintains clear boundaries between distinct analytical steps, whereas CoT conflates multiple operations within single reasoning thoughts. This structural distinction enables transparent, formally grounded reasoning that can be systematically verified and debugged at each stage.
To address these challenges, we introduce GateLens, a reasoning-enhanced LLM agent [miehling2025agentic] for reliable tabular data analysis to support decision-making in domain-specific contexts. We validate the architecture in the automotive software release domain, where the need for precise, transparent, and efficient analysis is particularly acute.
A key challenge in LLM-based analytical agents is the potential mismatch between natural-language reasoning and the actual computation implemented in generated code—a reasoning-to-code gap that becomes more pronounced as analytical complexity increases. GateLens addresses this challenge by integrating structured relational analysis with domain-specific expertise through a reasoning layer built on Relational Algebra (RA) as an extension of Chain-of-Thought (CoT). This reasoning layer systematically breaks down complex validation tasks into discrete, formally grounded analytical steps. We adopt this approach to address essential limitations of vanilla CoT for our work: its reasoning steps are opaque and often cannot be mapped directly to executable code, its operations are not compositional and cannot be independently reused or debugged, and its reasoning is unstructured with operations blended together without clear intent or formal grounding. In contrast, our RA-based reasoning layer aims to ensure that each step is independently interpretable and reusable, reasoning is grounded in formal relational algebra, and intent is made explicit through well-defined operations. Figure 1 illustrates this distinction between our structured RA-based approach and conventional CoT reasoning.
GateLens simplifies three critical aspects of release validation:
@@
  • We design an architecture optimized for time- and safety-critical environments, minimizing LLM invocations while preserving reasoning depth for reliable tabular analysis. GateLens operates in a zero-shot setting, avoiding the need for few-shot examples or multi-agent coordination, which improves generalizability, execution speed, reduces maintenance overhead, and enhances transparency.
@@ 2 Background and Motivation
This diversity is reflected not only in perspective but also in the level of granularity required from the underlying data. For example, senior management may ask high-level questions such as: “List all vehicles that have not yet received global approval,” seeking a consolidated overview without reference to schema details or subsystem-specific attributes. In contrast, release managers or verification engineers may pose highly specific queries such as: “What are the baseline, phase, RC, and EUF values for RM-320 in the latest test suite?” These questions rely on detailed knowledge of domain terminology and schema structure. Although both types of queries operate on the same data foundation, they demand substantially different views and levels of abstraction.
@@
These limitations highlight the need for enhanced solutions combining domain expertise with advanced analytical reasoning capabilities. Such solutions must facilitate efficient decision-making while maintaining high safety and quality standards in automotive software development [wang2024surveyl]. Effective intelligent assistants can transform the release decision process, enabling release managers to prioritize result interpretation and strategic decision-making over routine data analysis.
3 Approach and Methodology
@@ 3.2 Core Components
3.2.1 Query Interpreter
The query interpreter agent is responsible for converting user queries into formal RA expressions, providing a precise framework for analytical reasoning. Before initiating this translation, the agent consults the knowledge base, comprising the data schema and domain-specific context. The data schema provides a detailed understanding of the dataset, including its relational modeling, field descriptions, data types, and enriched metadata capturing domain-specific acronyms and terminology mappings. This glossary-enhanced schema is injected into the prompt context, enabling accurate resolution of domain-specific terms into formal schema attributes during RA construction.
To handle imprecise queries without compromising data privacy or exceeding LLM context windows, GateLens employs a selective strategy for exposing categorical information. For low-cardinality attributes (e.g., fields with only a few distinct categories such as test status), all valid options are explicitly enumerated within the schema metadata. In contrast, for high-cardinality attributes, actual database values are never exposed; instead, the schema specifies structural patterns or expected formats (e.g., standard prefixes or identifier conventions). This hybrid design keeps the prompt context compact and privacy-preserving while providing the LLM with sufficient context to generate accurate filtering conditions.
Using this information, the agent verifies whether the query is relevant and within the scope of the dataset [manik2021out]. This validation step ensures that only supported and meaningful queries are processed, improving both accuracy and efficiency. Once the query is confirmed to be in scope, the agent leverages the data schemas to interpret and decompose the query into formal RA expressions.
The agent’s primary function is to map natural language queries into formal RA expressions, enhancing LLM reasoning through structured decomposition [khot2022decomposed]. This approach extends traditional Chain-of-Thoughts (CoT) [wei2022chain] reasoning by constraining the model to think within a formal system framework [zhang2023igniting]. Instead of generating free-form solutions, the agent must express analytics using standard relational algebra notations through a limited set of standard operations: selection, projection, union, set difference, cartesian product, and rename as basic operations, as well as derived operations such as join, intersection, and division and complemented by aggregation functions like average, minimum, maximum, sum, and count.
@@ 3.2.2 Coder
3.3 Data Handling
A key architectural decision in GateLens is its indirect interaction with test data. Rather than exposing sensitive test data directly to LLM agents, which could raise privacy concerns [boudewijn2023privacy] and exceed input context limitations, the system operates on data schemas and relational models. This approach serves multiple critical purposes:
@@ 4. RQ4:
    To what extent does RA-based reasoning eliminate the need for in-context learning?
4.1 Experimental Setup
To address the research questions introduced in Section 4, we designed and conducted extensive experiments to evaluate the performance of GateLens.
@@ 4.1 Experimental Setup
The second benchmark is derived from real-world user queries collected from production logs at our partner company. These queries were sourced from the historical logs of an agentic system that employed a well-established tabular data reasoning approach combining CoT [wei2022chain] prompting with Self-Consistency (SC) [wang2022self]. This system was used in production to support software release analytics, and the collected queries reflect a wide range of user roles, query types, and domain-specific requirements. While this system performs effectively in many scenarios, its limitations become apparent as the range of roles and users expands, leading to a significant diversification of queries. This broader query diversity exposes the system’s reliance on few-shot examples, making it less capable of handling highly complex, ambiguous, or ill-defined queries that require greater flexibility and adaptability. Nevertheless, this preliminary system played a critical role in data collection for GateLens by providing query logs used to develop and validate our approach. From these logs, we filtered out near-duplicates and selected 244 frequently repeated unique queries, which we then organized into eight functional categories based on their purposes. The ground truth for the 244 real-world queries was established through a two-stage manual annotation process: two domain experts independently solved an initial subset of 20 queries to align on interpretation and expected outputs, after which the remaining queries were divided and annotated following the agreed-upon criteria, with periodic cross-checks to ensure consistency.
In order to assess GateLens’s performance, we run experiments with two large language models (LLMs): GPT-4o, a leading commercial model, and Llama 3.1 70b, a recently released open-source model. We also benchmark GateLens against the CoT+SC [khoee2024gonogo] agentic system currently used to support the company’s release decisions. Our comparative analysis is designed to quantify the improvements introduced by GateLens’s novel architecture.
To address the challenge of handling out-of-scope user queries during real-time interactions, GateLens incorporates an in-scope filtering mechanism as explained in Section 3.2. This mechanism ensures that the system only attempts to process queries that fall within its scope, thereby improving reliability and reducing errors. Performance evaluation focused on two key aspects:
@@
Based on these definitions, we calculate Precision, Recall, and F1 scores to assess the performance of the system. Precision ensures that incorrect results are minimized, recall ensures relevant queries are addressed, and the F1 score balances the two to provide an overall assessment of system performance. By relying on a closed-set benchmark with established ground truths, these metrics enable us to rigorously isolate and measure the impact of our architectural choices during lab validation, setting the stage for the real-world industrial evaluation detailed in Section LABEL:sec:lessons.
4.2 Performance in Addressing User Queries (RQ1)
Table 1: Comparison of average token consumption and reduction between CoT+SC and GateLens across four difficulty levels. The evaluation was conducted on the first benchmark, which consists of 50 designed queries with annotated difficulty levels, with both agents utilizing GPT-4o.
@@ 4.2 Performance in Addressing User Queries (RQ1)
Table 2: Performance comparison of GateLens and the CoT+SC system across different categories on the second benchmark, which consists of 244 real-world queries.
@@
We conducted experiments to compare the performance of GateLens across the two introduced benchmarks. The first benchmark, consisting of 50 queries categorized by difficulty levels, was used to evaluate and compare the performance of GateLens and CoT+SC. Both systems were tested using GPT-4o and Llama 3.1 70B as their underlying LLMs. The results are summarized in Table 1.
@@
For the second benchmark, results in Table 2 show that GateLens (GPT-4o) and CoT+SC (GPT-4o) significantly outperformed GateLens with Llama 3.1 70B. This performance disparity is primarily due to the strict code generation requirements of the task, including table filtering, merging strategies, and key-value mapping operations, where GPT-4o demonstrated markedly superior capabilities.
@@
Table 3: Comparison of average token consumption and reduction between CoT+SC and GateLens across four difficulty levels. The evaluation was conducted on the first benchmark, which consists of 50 designed queries with annotated difficulty levels, with both agents utilizing GPT-4o.
Level Agent Avg Input Tokens Avg Output Tokens Avg Total Tokens Token Reduction
1 CoT+SC 11,905 186 12,091
GateLens 2,239 420 2,658 $\downarrow$ 78%
2 CoT+SC 13,747 368 14,116
GateLens 2,428 701 3,129 $\downarrow$ 78%
3 CoT+SC 14,726 441 15,168
GateLens 2,432 698 3,130 $\downarrow$ 79%
4 CoT+SC 17,103 452 17,555
GateLens 2,505 847 3,352 $\downarrow$ 81%
In addition, compared to the CoT+SC solution, GateLens uses significantly fewer tokens due to its effective use of RA as the intermediate representation and its zero-shot architecture. In Table 3, we compare the token usage of both systems across different difficulty levels. Notably, while the CoT+SC approach requires increasingly massive input contexts, primarily to accommodate few-shot examples, GateLens maintains a highly compact input footprint. As the query difficulty increases, the token reduction becomes even more significant, demonstrating the efficiency and scalability of the RA-based approach in handling complex queries.
RQ1 findings: GateLens with GPT-4o achieved 100% F1 score on the first benchmark across all difficulty levels and 83.51% F1 score on the second benchmark with 244 real-world queries. It outperformed CoT+SC (GPT-4o) by approximately 13 percentage points on real-world queries, indicating that the RA reasoning mechanism effectively addresses both complex logical operations and real-world noise, such as typos and ambiguous field names.
4.3 Robustness: Handling Out of Scope and Imprecise Queries (RQ2)
@@
4.3.1 Handling Out of Scope Queries
We compared GateLens with other models; the results can be found in Table 4. The results demonstrate that GateLens with GPT-4o achieved the best performance, particularly in terms of precision, which is approximately 40% higher than other models, indicating GateLens’ ability to avoid generating incorrect results.
@@ 4.3.1 Handling Out of Scope Queries
Table 4: Model comparison for out-of-scope queries.
@@
4.3.2 Handling Imprecise Queries
@@
Table 5: Model comparison for imprecise queries.
@@
The results of these experiments are presented in Table 5. As shown, GateLens with GPT-4o demonstrates the best overall performance. In terms of precision, all methods performed relatively well, indicating that when results are generated, they are likely to be correct. However, our method significantly outperformed the others in recall, highlighting its ability to handle a larger portion of the imprecise queries. As a result, GateLens with GPT-4 achieved a substantially higher F1 score compared to other methods, demonstrating that it not only processes most queries but also produces accurate results for them.
@@
RQ2 findings: GateLens demonstrates high robustness in real-world conditions. It effectively filters out-of-scope inputs, achieving approximately 40% higher precision than the baseline system. Furthermore, it handles imprecise or informal queries with superior performance, more than doubling the recall (78% vs. 36%). This confirms that the system can manage ambiguous user inputs without sacrificing the accuracy of the generated code.
4.4 Effectiveness of the RA module (RQ3)